What is the measure of the third sides of a triangle, that is already measured with 5meters and 4meters?
Answer by oberobic(2304)

There is no answer to this question because the triangle could any shape.
If the angle between the two known sides was 1 degree, then the third side would relatively short.
If the angle between the two known sides was 179 degrees, then the third side would be relatively long.
.
If you have studied the law of cosines, you know that it extends the Pythagorean Theorem to any angle.
c^2 = a^2 + b^2 - 2*a*b*cos(x), where x is the angle between the known sides a and b.
.
(I used 'x' where most texts would say the angle was a Greek letter.
But I don't have a Greek letter to type into this reply. Sorry.)
.
To use the law of cosines, you have to know the measure of the angle.
.
cos 1° = 0.99984769515639
.
c^2 = 4^2 + 5^2 - 2*(4)*(5)*(0.99984769515639)
.
c^2 = 1.00609219
.
c = sqrt(1.00609219) = 1.00304147
.
VS.
.
cos 179° = -0.99984769515639
.
c^2 = 4^2 + 5^2 - 2*(4)*(5)*(-0.99984769515639)
.
c^2 = 80.9939078
.
c = sqrt(80.9939078) = 8.99966154
.
Done.